OpenStudy (anonymous):

g(x) = x / (x-2) g'(x) = ?

OpenStudy (anonymous):

Using quotient rule you have: \[f'(x)=\frac{(1)(x-2)-(x)(1)}{(x-2)^2}=\frac{-2}{(x-2)^2}\]

OpenStudy (anonymous):

you forgot to put g'(x) instead of f'(x)...now Bovice is confused...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Lol. it should be g'(x)

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

oh yeah totally confused :s lol